# DaWin: Training-free Dynamic Weight Interpolation for Robust Adaptation

### v3

@@ Abstract
Adapting a pre-trained foundation model on downstream tasks should ensure robustness against distribution shifts without the need to retrain the whole model. Although existing weight interpolation methods are simple yet effective, we argue their static nature limits downstream performance while achieving efficiency. In this work, we propose DaWin, a training-free dynamic weight interpolation method that leverages the entropy of individual models over each unlabeled test sample to assess model expertise, and compute per-sample interpolation coefficients dynamically. Unlike previous works that typically rely on additional training to learn such coefficients, our approach requires no training. Then, we propose a mixture modeling approach that greatly reduces inference overhead raised by dynamic interpolation. We validate DaWin on the large-scale visual recognition benchmarks, spanning 14 tasks across robust fine-tuning – ImageNet and derived five distribution shift benchmarks – and multi-task learning with eight classification tasks. Results demonstrate that DaWin achieves significant performance gain in considered settings, with minimal computational overhead. We further discuss DaWin’s analytic behavior to explain its empirical success. Here is our code.
@@ 2.1 Background
Classification under distribution shift. We consider a classification problem over a domain $\mathcal{X}$ with input $x$ and a label space $\mathcal{Y}=\{1,...,C\}$, where the goal is to approximate the true labeling function $h:\mathcal{X}\rightarrow\mathcal{Y}$ with a parametric model $f:\mathcal{X}\rightarrow\Delta^{C-1}$. This model maps $x$ to $C-1$ dimensional simplex, which aims to minimize error $l:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ on inputs drawn from any potential target distribution $P_{T}$. In the robust fine-tuning scenario (Wortsman et al., 2022b; Kumar et al., 2022), the target distribution $P_{T}$, where a pre-trained model $f$ being adapted to, is typically assumed to be covariate-shifted version (OOD) of the ID source distribution $P_{S}$. Both distributions share the same class label space $\mathcal{Y}$, and have the same conditional distribution over target labels, i.e., $P_{S}(y|x)=P_{T}(y|x)$, but have different marginal distributions over input $P_{S}(x)\neq P_{T}(x)$.
Model merging. Let $f(\cdot;\theta_{0})$ and $f(\cdot;\theta_{1})$ denote models that are individually trained on the same or different datasets but have identical architecture. For example, $f(\cdot;\theta_{0})$ could represent a pre-trained model such as CLIP (Radford et al., 2021), while $f(\cdot;\theta_{1})$ is the fine-tuned counterpart on a particular downstream task. Model merging approach (Wortsman et al., 2022b; a) constructs a merged model $f(\cdot;\theta_{\lambda})$, which achieves a better trade-off between ID and OOD performance than the individual models by interpolating in the weight space $\theta_{\lambda}=(1-\lambda)\theta_{0}+\lambda\theta_{1}$. We will use the terms interpolation and merging interchangeably. Throughout the paper, we use the term static interpolation to denote methods that induce a single merged model corresponding to a single interpolation coefficient applied for all test samples. In contrast, dynamic interpolation refers to methods that yield multiple merged models, with interpolation coefficients depending on the sample $x$ or domain $\mathcal{X}$.
@@ 2.2 Pilot Study
Our goal. We begin by conducting a pilot study to understand the benefits of dynamic interpolation, which adapts interpolation coefficients at a finer granularity (such as sample level, $\lambda(x)$), as opposed to using global coefficient $\lambda$ for all samples (Wortsman et al., 2022b; a). To explore the upper limits of these methods, we experiment with ground truth labels as an oracle [^1] We use the term oracle to denote a scenario where we utilize the ground truth test label, and it should not be confused with optimal. Optimality of $\lambda^{*}(x)$ regarding weight interpolation is further discussed in Tab C. to estimate upper-bound performance and understand the maximum potential of model interpolation methods. This study will further guide our method design in Sec. 3 by approximating the oracle performance.
@@
In this experiment, we hypothesize that (1) Fine-grain interpolation, which adapts coefficients at a finer level (such as the sample level), can lead to substantial improvements compared to static interpolation, such as WiSE-FT. (2) Per-sample interpolation coefficients can effectively estimated with X-entropy-based model expertise measurement.
@@ 3.1 Revisiting Entropy as a Measure of Model Expertise
While the pilot study in the previous section provides encouraging results, those oracle methods cannot work in practice. This is because we do not have access to the ground truth labels for incoming test samples. This leads us to the question: how can we reliably estimate the interpolation coefficient solely based on the test input $x$? There is extensive literature which adopts entropy as a proxy of X-entropy (Grandvalet & Bengio, 2004; Chapelle & Zien, 2005; Shu et al., 2018; Wang et al., 2021; Prabhu et al., 2021). The rationale behind using entropy as a surrogate is grounded in the observation that the entropy of a sample’s prediction distribution strongly correlates with X-entropy, even under distribution shifts (Wang et al., 2021) those models have not explicitly trained on. This work presents the first attempt to leverage the sample-wise entropy to measure each model’s expertise to determine the interpolation coefficients given each test sample. It is worth noting that this approach differs from AdaMerging (Yang et al., 2024b), which seeks to learn a global coefficient inducing a single merged model to minimize the expected entropy over the entire test set.
@@
Here, aside from the FalseFalse case where both models failed to produce reliable predictions [^3] However, even individuals fail, an interpolated model can yield correct predictions (Yong et al., 2024)., we observe strong correlations (Schober et al., 2018) across the remaining splits and the entire dataset. This implies that entropy is a reasonable proxy to the X-entropy for estimating per-sample model expertise and, ultimately, computing sample-wise interpolation coefficients.
@@ 3.2 Dynamic Interpolation via Entropy Ratio
Here, $H$ represents the entropy over the output probability distribution of model $f(\cdot;\theta)$ given input $x$. The value of $\lambda(x)$ approaches 1 when model $f(\cdot;\theta_{1})$ exhibits lower entropy (greater expertise), whereas it approaches 0 if the entropy of model $f(\cdot;\theta_{1})$ becomes higher. Figure 3 illustrates the overall framework. Intuitively, this approach is analogous to classifier selection methods (Giacinto & Roli, 2001; Ko et al., 2008), which aims to dynamically select suitable classifier(s) given a test sample. However, DaWin differs in terms of its motivation, expertise metric, and goal as we pursue finding the interpolation coefficients that may induce better performance compared with the best model selection method (See Table 5). We further discuss the connection between DaWin and the selection-based method in Sec. 6. Note that our sample-wise expertise estimations do not require any hyperparameters for computing interpolation coefficients. This eliminates the need for intensive hyperparameter tuning often required in prior works (Wortsman et al., 2022b; Ilharco et al., 2023) to achieve the best result. Meanwhile, if we assume access to all test samples simultaneously, we can refine some unstable coefficient estimations by introducing per-domain expertise offset. We use this offset adjustment by default (Sec. A.1). Besides, to bypass using overconfident output (Guo et al., 2017) during expertise estimation, we conduct the temperature scaling on ID validation set.
Sample-wise entropy valley hypothesis. One of the most popular hypotheses that explain the success of weight interpolations is linear mode connectivity (LMC; Frankle et al. (2020)), which ensures that interpolated models are also laid on good solution space as well as low-loss converged individual models. Although entropy is a concave function for the probability density, we believe that a similar statement can be claimed, which we refer to sample-wise entropy valley as follows:
@@ 4.2 Results on Robust Fine-tuning
Better performance trade-off between ID and OOD. In Figure 4, we present the accuracy for ID ($x$-axis) and OOD average ($y$-axis) across three backbone models, along with baseline performance. Achieving superior ID/OOD trade-offs beyond or even comparable to the WiSE-FT hyperparameter sweep is challenging. Nevertheless, DaWin provides a remarkably better performance trade-off compared to baseline methods. It is also worth noting that DaWin does not require any hyperparameter tuning. Instead, it dynamically generates interpolation coefficients solely based on the trained models’ output. This distinguishes DaWin from other hyperparameter-intensive fine-tuning such as CaRot (Oh et al., 2024) and Lipsum-FT (Nam et al., 2024) or existing interpolation methods (Wortsman et al., 2022b). Table 2 provides detailed train/inference costs and ID/OOD accuracies. Compared with Model Soups (Uniform and Greedy Soups) and Model Stock, which require more than two models to ensure diversity among interpolation candidate models (Wortsman et al., 2022a) or periodic interpolations during training (Jang et al., 2024), DaWin operates with only a single fine-tuned model while achieving significantly better performance trade-off. These trends hold across other backbones (see Tables 4.2 and 4.2), verifying DaWin’s generality in terms of modeling scale. Although DaWin compromises the inference-time efficiency for accuracy, given that existing works typically require hyperparameter tuning in practice (Wortsman et al., 2022b), relative computation overhead is minor given that we set $(K+M)\lesssim H$ in all our experiments (See Figure 8).
@@ 4.3 Results on Multi-task Learning
We now evaluate existing merging approaches and our DaWin on multi-task learning benchmarks with CLIP ViT-B/32. Following the standard evaluation protocol (Yang et al., 2024b), we have $M$ tasks and models individually fine-tuned on each task (where $M=8$). We do not know where each test sample arises from during evaluation and cannot choose true experts per domain. While both AdaMerging and DaWin use unlabeled testset, DaWin produces dynamic merged models given tasks or samples, whereas AdaMerging induces a single merged model for all tasks by default.
@@ 5 Related Work
Robust fine-tuning aims to adapt a model on a target task while preserving the generalization capability learned during pre-training. A straightforward approach injects regularization into the learning objective. For example, Ju et al. (2022) proposed a regularization motivated by Hessian analysis, Tian et al. (2023a; b) devised a trainable projection method to constrain the parameter space, CAR-FT (Mao et al., 2024) devised the context-awareness regularization, and CaRot (Oh et al., 2024) introduced a regularization based on singular values. Another line of works modifies the training procedure to keep the pre-trained knowledge by decoupling the tuning of a linear head from the entire model (Kumar et al., 2022), employing a data-dependent tunable module (Lee et al., 2023), utilizing bi-level optimization (Choi et al., 2024), or mimicking the pre-training procedure (Goyal et al., 2023). In contrast, weight interpolation approaches emerged as an effective yet efficient solution that conducts simple interpolation of individual model weights (Izmailov et al., 2018; Wortsman et al., 2022a; b; Jang et al., 2024). Unlike existing works inducing a single interpolated model, we propose a dynamic interpolation method that produces per-sample models for better adaptation.
@@ 6 Discussion
Accuracy and runtime trade-off. While we focus on improving accuracy through model merging approaches, it is crucial to ensure the extra computation demands of DaWin during inference are manageable. Fig. 8 presents trade-offs between the accuracy and wall-clock time for various merging methods. For methods requiring hyperparameter tuning, e.g., WiSE-FT and Task Arithmetic, the wall-block time (logarithm of second) reflects a cumulative time for evaluations across all hyperparameters. For methods like DaWin or AdaMerging, we include the time required for additional workloads. In both settings, DaWin shows favorable trade-offs that outperform the most efficient one in terms of accuracy while its runtime is far less than computation-heavier methods such as per sample interpolation without mixture modeling (Sample-wise Interp.) or AdaMerging. Thus, DaWin provides a high-performing merging solution that can be flexibly adapted considering a given cost budget.
@@ A.3 Additional empirical results
Figure F: Entropy comparison on ImageNet distribution shiftㄴ with CLIP ViT-B/32. Across IN, IN-V2, IN-R, IN-A, IN-S, and ObjNet from top to bottom, we visualize the final output entropy from each model on the entire dataset and four different splits based on the correctness of zero-shot (ZS) and fine-tuned (FT) models. Compared with individual models, weight averaging (WA) induces lower entropy overall, and our DaWin achieves the lowest entropy across all splits.
@@
Similar to the linear mode connectivity phenomenon that prevails between pre-trained and fine-tuned checkpoints of foundation models (Wortsman et al., 2022b; Ramé et al., 2023) which indicates the interpolations between two checkpoints do not increase the expected loss (e.g., X-entropy) on downstream, in Section 3.2, we hypothesized the expected entropy of model output may also show similar convexity trend over interpolations between pre-trained and fine-tuned model checkpoints. In Figure H, we see that this claim empirically holds for the six datasets including ID (ImageNet) and OOD (its five variances), i.e., there exists an interpolated model that achieves smaller entropy than individual models (ZS and FT). Besides, we further observed that there exists a per-sample dynamic interpolation method that achieves much smaller entropy compared to the static interpolation method. Therefore, our two statements in the sample-wise entropy valley hypothesis show empirical evidence.
@@ A.5 Limitation and future work
Meanwhile, the increased amount of computation during inference time is another limitation of DaWin. However, we note that scaling test-time computation can be a cost-effective solution for foundation models to address challenging tasks and data in the wild (Snell et al., 2024; Jaech et al., 2024; Muennighoff et al., 2025). Given the observations that modern multimodal foundation models still struggle with out-of-distribution query (OpenAI, 2024; Zhang et al., 2024; Oh et al., 2025), improving the robustness of foundation models by allocating more budgets for test-time computation through dynamic merging, such as DaWin, is worth to be investigated further.

### v4

@@ Abstract
Adapting a pre-trained foundation model on downstream tasks should ensure robustness against distribution shifts without the need to retrain the whole model. Although existing weight interpolation methods are simple yet effective, we argue that their static nature limits downstream performance while achieving efficiency. In this work, we propose DaWin, a training-free dynamic weight interpolation method that leverages the entropy of individual models over each unlabeled test sample to assess model expertise, and compute per-sample interpolation coefficients dynamically. Unlike previous works that typically rely on additional training to learn such coefficients, our approach requires no training. Then, we propose a mixture modeling approach that greatly reduces inference overhead raised by dynamic interpolation. We validate DaWin on the large-scale visual recognition benchmarks, spanning 14 tasks across robust fine-tuning – ImageNet and derived five distribution shift benchmarks – and multi-task learning with eight classification tasks. Results demonstrate that DaWin achieves significant performance gain in considered settings, with minimal computational overhead. We further discuss DaWin’s analytic behavior to explain its empirical success. Here is our code.
@@ 2.1 Background
Classification under distribution shift. We consider a classification problem over a domain $\mathcal{X}$ with input $x$ and a label space $\mathcal{Y}=\{1,...,C\}$, where the goal is to approximate the true labeling function $h:\mathcal{X}\rightarrow\mathcal{Y}$ with a parametric model $f:\mathcal{X}\rightarrow\Delta^{C-1}$. This model maps $x$ to $C-1$ dimensional simplex, which aims to minimize error $l:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ on inputs drawn from any potential target distribution $P_{T}$. In the robust fine-tuning scenario (Wortsman et al., 2022b; Kumar et al., 2022), the target distribution $P_{T}$, where a pre-trained model $f$ being adapted to, is typically assumed to be a covariate-shifted version (OOD) of the ID source distribution $P_{S}$. Both distributions share the same class label space $\mathcal{Y}$, and have the same conditional distribution over target labels, i.e., $P_{S}(y|x)=P_{T}(y|x)$, but have different marginal distributions over input $P_{S}(x)\neq P_{T}(x)$.
Model merging. Let $f(\cdot;\theta_{0})$ and $f(\cdot;\theta_{1})$ denote models that are individually trained on the same or different datasets but have identical architecture. For example, $f(\cdot;\theta_{0})$ could represent a pre-trained model such as CLIP (Radford et al., 2021), while $f(\cdot;\theta_{1})$ is the fine-tuned counterpart on a particular downstream task. Model merging approach (Wortsman et al., 2022b; a) constructs a merged model $f(\cdot;\theta_{\lambda})$, which achieves a better trade-off between ID and OOD performance than the individual models by interpolating in the weight space $\theta_{\lambda}=(1-\lambda)\theta_{0}+\lambda\theta_{1}$. We will use the terms interpolation and merging interchangeably. Throughout the paper, we use the term static interpolation to denote methods that induce a single merged model corresponding to a single interpolation coefficient applied to all test samples. In contrast, dynamic interpolation refers to methods that yield multiple merged models, with interpolation coefficients depending on the sample $x$ or domain $\mathcal{X}$.
@@ 2.2 Pilot Study
Our goal. We begin by conducting a pilot study to understand the benefits of dynamic interpolation, which adapts interpolation coefficients at a finer granularity (such as sample level, $\lambda(x)$), as opposed to using a global coefficient $\lambda$ for all samples (Wortsman et al., 2022b; a). To explore the upper limits of these methods, we experiment with ground truth labels as an oracle [^1] We use the term oracle to denote a scenario where we utilize the ground truth test label, and it should not be confused with optimal. Optimality of $\lambda^{*}(x)$ regarding weight interpolation is further discussed in Tab C. to estimate upper-bound performance and understand the maximum potential of model interpolation methods. This study will further guide our method design in Sec. 3 by approximating the oracle performance.
@@
In this experiment, we hypothesize that (1) Fine-grain interpolation, which adapts coefficients at a finer level (such as the sample level), can lead to substantial improvements compared to static interpolation, such as WiSE-FT. (2) Per-sample interpolation coefficients can be effectively estimated with X-entropy-based model expertise measurement.
@@ 3.1 Revisiting Entropy as a Measure of Model Expertise
While the pilot study in the previous section provides encouraging results, those oracle methods cannot work in practice. This is because we do not have access to the ground truth labels for incoming test samples. This leads us to the question: how can we reliably estimate the interpolation coefficient solely based on the test input $x$? There is extensive literature which adopts entropy as a proxy of X-entropy (Grandvalet & Bengio, 2004; Chapelle & Zien, 2005; Shu et al., 2018; Wang et al., 2021; Prabhu et al., 2021). The rationale behind using entropy as a surrogate is grounded in the observation that the entropy of a sample’s prediction distribution strongly correlates with X-entropy, even under distribution shifts (Wang et al., 2021) those models have not been explicitly trained on. This work presents the first attempt to leverage the sample-wise entropy to measure each model’s expertise to determine the interpolation coefficients given each test sample. It is worth noting that this approach differs from AdaMerging (Yang et al., 2024b), which seeks to learn a global coefficient inducing a single merged model to minimize the expected entropy over the entire test set.
@@
Here, aside from the FalseFalse case where both models failed to produce reliable predictions [^3] However, even if individuals fail, an interpolated model can yield correct predictions (Yong et al., 2024)., we observe strong correlations (Schober et al., 2018) across the remaining splits and the entire dataset. This implies that entropy is a reasonable proxy to the X-entropy for estimating per-sample model expertise and, ultimately, computing sample-wise interpolation coefficients.
@@ 3.2 Dynamic Interpolation via Entropy Ratio
Here, $H$ represents the entropy over the output probability distribution of model $f(\cdot;\theta)$ given input $x$. The value of $\lambda(x)$ approaches 1 when model $f(\cdot;\theta_{1})$ exhibits lower entropy (greater expertise), whereas it approaches 0 if the entropy of model $f(\cdot;\theta_{1})$ becomes higher. Figure 3 illustrates the overall framework. Intuitively, this approach is analogous to classifier selection methods (Giacinto & Roli, 2001; Ko et al., 2008), which aims to dynamically select suitable classifier(s) given a test sample. However, DaWin differs in terms of its motivation, expertise metric, and goal as we pursue finding the interpolation coefficients that may induce better performance compared with the best model selection method (See Table 5). We further discuss the connection between DaWin and the selection-based method in Sec. 6. Note that our sample-wise expertise estimations do not require any hyperparameters for computing interpolation coefficients. This eliminates the need for intensive hyperparameter tuning often required in prior works (Wortsman et al., 2022b; Ilharco et al., 2023) to achieve the best result. Meanwhile, if we assume access to all test samples simultaneously, we can refine some unstable coefficient estimations by introducing per-domain expertise offsets. We use this offset adjustment by default (Sec. A.1). Besides, to bypass using overconfident output (Guo et al., 2017) during expertise estimation, we conduct the temperature scaling on ID validation set.
Sample-wise entropy valley hypothesis. One of the most popular hypotheses that explains the success of weight interpolations is linear mode connectivity (LMC; Frankle et al. (2020)), which ensures that interpolated models are also laid on good solution space as well as low-loss converged individual models. Although entropy is a concave function for the probability density, we believe that a similar statement can be claimed, which we refer to sample-wise entropy valley as follows:
@@ 4.2 Results on Robust Fine-tuning
Better performance trade-off between ID and OOD. In Figure 4, we present the accuracy for ID ($x$-axis) and OOD average ($y$-axis) across three backbone models, along with baseline performance. Achieving superior ID/OOD trade-offs beyond or even comparable to the WiSE-FT hyperparameter sweep is challenging. Nevertheless, DaWin provides a remarkably better performance trade-off compared to baseline methods. It is also worth noting that DaWin does not require any hyperparameter tuning. Instead, it dynamically generates interpolation coefficients solely based on the trained models’ output. This distinguishes DaWin from other hyperparameter-intensive fine-tuning, such as CaRot (Oh et al., 2024) and Lipsum-FT (Nam et al., 2024), or existing interpolation methods (Wortsman et al., 2022b). Table 2 provides detailed train/inference costs and ID/OOD accuracies. Compared with Model Soups (Uniform and Greedy Soups) and Model Stock, which require more than two models to ensure diversity among interpolation candidate models (Wortsman et al., 2022a) or periodic interpolations during training (Jang et al., 2024), DaWin operates with only a single fine-tuned model while achieving significantly better performance trade-off. These trends hold across other backbones (see Tables 4.2 and 4.2), verifying DaWin’s generality in terms of modeling scale. Although DaWin compromises the inference-time efficiency for accuracy, given that existing works typically require hyperparameter tuning in practice (Wortsman et al., 2022b), relative computation overhead is minor given that we set $(K+M)\lesssim H$ in all our experiments (See Figure 8).
@@ 4.3 Results on Multi-task Learning
We now evaluate existing merging approaches and our DaWin on multi-task learning benchmarks with CLIP ViT-B/32. Following the standard evaluation protocol (Yang et al., 2024b), we have $M$ tasks and models individually fine-tuned on each task (where $M=8$). We do not know where each test sample arises from during evaluation, and cannot choose true experts per domain. While both AdaMerging and DaWin use an unlabeled testset, DaWin produces dynamic merged models given tasks or samples, whereas AdaMerging induces a single merged model for all tasks by default.
@@ 5 Related Work
Robust fine-tuning aims to adapt a model on a target task while preserving the generalization capability learned during pre-training. A straightforward approach injects regularization into the learning objective. For example, Ju et al. (2022) proposed a regularization motivated by Hessian analysis, Tian et al. (2023a; b) devised a trainable projection method to constrain the parameter space, CAR-FT (Mao et al., 2024) devised the context-awareness regularization, and CaRot (Oh et al., 2024) introduced a regularization based on singular values. Another line of work modifies the training procedure to keep the pre-trained knowledge by decoupling the tuning of a linear head from the entire model (Kumar et al., 2022), employing a data-dependent tunable module (Lee et al., 2023), utilizing bi-level optimization (Choi et al., 2024), or mimicking the pre-training procedure (Goyal et al., 2023). In contrast, weight interpolation approaches emerged as an effective yet efficient solution that conducts simple interpolation of individual model weights (Izmailov et al., 2018; Wortsman et al., 2022a; b; Jang et al., 2024). Unlike existing works inducing a single interpolated model, we propose a dynamic interpolation method that produces per-sample models for better adaptation.
@@ 6 Discussion
Accuracy and runtime trade-off. While we focus on improving accuracy through model merging approaches, it is crucial to ensure the extra computation demands of DaWin during inference are manageable. Fig. 8 presents trade-offs between the accuracy and wall-clock time for various merging methods. For methods requiring hyperparameter tuning, e.g., WiSE-FT and Task Arithmetic, the wall-clock time (logarithm of second) reflects a cumulative time for evaluations across all hyperparameters. For methods like DaWin or AdaMerging, we include the time required for additional workloads. In both settings, DaWin shows favorable trade-offs that outperform the most efficient one in terms of accuracy while its runtime is far less than computation-heavy methods such as per-sample interpolation without mixture modeling (Sample-wise Interp.) or AdaMerging. Thus, DaWin provides a high-performing merging solution that can be flexibly adapted considering a given cost budget.
@@ A.3 Additional empirical results
Figure F: Entropy comparison on ImageNet distribution shifts with CLIP ViT-B/32. Across IN, IN-V2, IN-R, IN-A, IN-S, and ObjNet from top to bottom, we visualize the final output entropy from each model on the entire dataset and four different splits based on the correctness of zero-shot (ZS) and fine-tuned (FT) models. Compared with individual models, weight averaging (WA) induces lower entropy overall, and our DaWin achieves the lowest entropy across all splits.
@@
Similar to the linear mode connectivity phenomenon that prevails between pre-trained and fine-tuned checkpoints of foundation models (Wortsman et al., 2022b; rame2023model) which indicates the interpolations between two checkpoints do not increase the expected loss (e.g., X-entropy) on downstream, in Section 3.2, we hypothesized the expected entropy of model output may also show similar convexity trend over interpolations between pre-trained and fine-tuned model checkpoints. In Figure H, we see that this claim empirically holds for the six datasets including ID (ImageNet) and OOD (its five variances), i.e., there exists an interpolated model that achieves smaller entropy than individual models (ZS and FT). Besides, we further observed that there exists a per-sample dynamic interpolation method that achieves much smaller entropy compared to the static interpolation method. Therefore, our two statements in the sample-wise entropy valley hypothesis show empirical evidence.
@@ A.5 Limitation and future work
Meanwhile, the increased amount of computation during inference time is another limitation of DaWin. However, we note that scaling test-time computation can be a cost-effective solution for foundation models to address challenging tasks and data in the wild (snell2024scaling; jaech2024openai; muennighoff2025s1). Given the observations that modern multimodal foundation models still struggle with out-of-distribution query (openai2024gpt4vsystemcard; zhang2024out; oh2025understanding), improving the robustness of foundation models by allocating more budgets for test-time computation through dynamic merging, such as DaWin, is worth to be investigated further.